OpenStudy (anonymous):

Tony worked as a substitute teacher last month. He substituted a total of 18 days. The first part of the month he worked in tucson, Arizona, where he was paid $75 a day. Then he moved to Oregon, where he was paid $135 a day to substitute. During the month he earned $2010. How many days did he substitute in Tucson? How many days did he substitute in Oregon?

OpenStudy (anonymous):

Deja vu....

OpenStudy (anonymous):

x+y=18 75x+135y=2010 y=18-x